OpenStudy (anonymous):

Jose has a number of nickels and dimes in his bank. Their total value is less than or equal to $10. He wrote a linear inequality in which x represents the number of nickels and y represents the number of dimes. Use the graphing tool to graph his equation. 0.05x + 0.10y ≤ 10 Which shows a possible number of nickels and dimes that Jose can have? a. 40 nickels and 50 dimes b. 60 nickels and 80 dimes c. 120 nickels and 50 dimes d. 150 nickels and 30 dimes

OpenStudy (anonymous):

@mathmath333 Anything ?

OpenStudy (mathmath333):

u have to put all options and see which one is/are correct for example a. 40 nickels and 50 dimes \(\large\tt \begin{align} \color{black}{0.05x + 0.10y ≤ 10\\ 0.05x + 0.10y \\ =0.05\times 40 + 0.10\times 50\\ =\dfrac{5}{100}\times 40 +\dfrac{10}{100}\times 50\\ =2 +5\\~\\ \Large =7≤10}\end{align}\) so first option is correct like these try for other options

OpenStudy (anonymous):

Thank you

OpenStudy (mathmath333):

it will very easy when u plot like this in graph http://www.wolframalpha.com/input/?i=plot+0.05x+%2B+0.10y+%E2%89%A4+10